OpenStudy (anonymous):

How many solutions does this system have? -3x+6y=10 -3x+6y=-4 A. None B. One C. Infinite D. Two

OpenStudy (solomonzelman):

subtract the equations from each other. -3x-(-3x) +6y-(+6y) = 10 - 4 -3x+3x +6y-6y = 10 - (-4) 0x + 0y = 14 0=14 so therefore it has no solution, see?

OpenStudy (solomonzelman):

I wrote 6 instead of 14, sorry.

OpenStudy (anonymous):

That makes sense.

OpenStudy (solomonzelman):

K